2. Regression analysis was applied between sales (in $1,000) and advertising (in $100), and the following regression function was obtained
Y=80+6.2x
Based on the above estimated regression line, if advertising is $10,000, then the point estimate for sales (in dollars) is

According to my quiz the answer is 700,000 , please help me figure out how.
Answer #1

Since the equation is Y= 80+6.2 *X where X is the advertising cost, and since the equation which is given was calculated from the sample of which advertising was in $100. So the advertising value will be $100k here.

So, now the sales is calculated as:

Y= 80+6.2*100 = $700k.

Now since the sales is given in $1000 hence the correct sale will be 700*1000 = 700,100.